# Critical modular lattices in the Gaussian core model

### v1

@@ Abstract.
We discuss the local analysis of Gaussian potential energy of modular lattices. We show for instance that the $3$-modular $12$-dimensional Coxeter-Todd lattice and the $2$-modular $16$-dimensional Barnes-Wall lattice, which both provide excellent sphere packings, are not, even locally, universally optimal (in the sense of Cohn and Kumar).
@@ 1.1. The Gaussian core model
for $\alpha>0$. Point configurations which interact via Gaussian potential functions $r\mapsto e^{-\alpha r^{2}}$ are referred to as the Gaussian core model. They are natural physical systems (see Stillinger [29]) and they are mathematically quite general. This embeds into the larger context of universal optimality: Following Cohn and Kumar [4], a point configuration is called universally optimal if it minimizes potential energy among all configurations of the same point density for every completely monotonic function of squared distance. It is an important observation that after fixing a point density, a point configuration is universally optimal if and only if it minimizes Gaussian potential energy for all $\alpha>0$. This is a consequence of a theorem of Bernstein (see Widder [34, Theorem 12b, page 161]). See Cohn, Kumar, Miller, Radchenko and Viazovska [6] for further discussion.
In this seminal paper [6] the authors proved that the $E_{8}$ root lattice in dimension $8$ and the Leech lattice $\Lambda_{24}$ in dimension $24$ are universally optimal point configurations in their respective dimensions. Further, the authors proved that these lattices are the unique minimizers among all periodic point configurations.
@@
One of the main goals of [14], motivated by Regev and Stephens-Davidowitz [25], was to find a concrete example of an even unimodular lattice that is a local maximum for some energy parameter $\alpha$ and the associated Gaussian potential. The first such example in this class of lattices can be found dimension $32$, every extremal even unimodular lattice in dimension $32$ is a local maximum for certain values of $\alpha$, a rigorous proof of this statement was worked out specifically for $\alpha=\pi$. It was proved that $E_{8}\perp E_{8}$ in dimension $16$ and every Niemeier lattice with a decomposable root system in dimension $24$ are saddle points if $\alpha$ is large enough. While all even unimodular lattices in dimensions up to $24$ turned out to be critical points in the Gaussian core model, there are non-critical even unimodular lattices in dimension $32$.
The aim of this paper is to extend the local analysis to a larger class of lattices, including dimensions in which no even unimodular lattice can exist. For this we consider the class of (extremal) modular lattices. This is in part motivated by an observation of Sarnak and Strömbergsson [26]: A lattice (of point density $1$) cannot be universally optimal if the vector lengths in $L$ differ from that of its dual lattice. It seems noteworthy to mention that $A_{2}$ and $D_{4}$, which are locally universally optimal among lattices, are actually $2$-modular lattices, that is $A_{2}\cong\sqrt{2}A_{2}^{\#}$ and $D_{4}\cong\sqrt{2}D_{4}^{\#}$. From the general definition of an $\ell$-modular lattice below, it becomes clear that $\ell$-modular lattices satisfy this necessary condition and therefore seem good candidates to investigate if one is interested in universal optimality.
@@ 1.2. Modular lattices in the Gaussian core model
Modular lattices generalize unimodular lattices. This generalization is closely related to the theta series of a lattice $L$, which is a function of a variable $\tau$ in the upper half plane given by
@@
We now move to a larger class of lattices, for which $\Theta_{L}$ has similarly nice properties. Let $\ell>0$. A similarity $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of norm $\ell$ is a linear map satisfying
@@
1.3. Structure of the paper and main results
@@
The application of this strategy requires bounds on the coefficients of modular forms, which we obtain separately for Eisenstein series and cusp forms. This is the content of Section 4. Here we collect explicit bounds for the coefficients of cusp forms for $\Gamma_{0}(2)$ and $\Gamma_{0}(3)$, which are available in the literature.
@@
We can summarize these results as follows.
Theorem 1.2.
The $3$-modular $12$-dimensional Coxeter-Todd lattice $K_{12}$ and the $2$-modular $16$-dimensional Barnes-Wall lattice $\operatorname{BW}_{16}$ are not locally universally optimal.
In fact for both lattices we find that there exists some specific $\alpha$ such that these lattices are locally maximal for $\mathcal{E}(\alpha,L)$, this is the subject of in Propositions 5.1 and 5.3. Note that by combining the computations there with Theorem 3.7 one obtains that both lattices are local minima for $\mathcal{E}(\alpha,L)$ if $\alpha$ is large. This seems plausible as both lattices are the best known sphere packings in their respective dimension and the energy minimization problem converges to the sphere packing problem for $\alpha\rightarrow\infty$. These lattices posses a lot of structure that simplifies the computation of the eigenvalues of the Hessian of $\mathcal{E}(\alpha,L)$ significantly; this comes from the fact that the shells of these lattices form spherical $4$-designs. Recall that if the first shell, i.e. the shell associated to minimal vectors of the lattice, forms a spherical $4$-design, then the lattice is called strongly perfect, a notion by Venkov [31]. It is known that strongly perfect lattices are extreme in the geometric sense, i.e. they are strict local maxima of the sphere packing density function for lattices. By a classical result of Voronoi [32] a lattice is extreme if and only if it is perfect and eutactic. Here a lattice $L$ is perfect if the span of the rank one matrices $xx^{\sf{T}}$, where $x$ runs through the set of minimal vectors of $L$, is equal to the space of symmetric matrices; $L$ is eutactic if the identity matrix is in the interior of the convex cone spanned by the rank one matrices $xx^{\sf{T}}$ from before.
@@
If $L$ is an even $\ell$-modular lattice we can compute the eigenvalues of the Hessian of energy at $L$ under some assumptions on a related space of modular forms, this is the content of Theorem 3.3. Building on this, if $L$ is an extremal even $\ell$-modular lattice, where $\ell$ is prime and $1+\ell$ divides $24$, we can determine the asymptotic behavior of $L$ in the Gaussian core model, as $\alpha\rightarrow\infty$, under the assumptions made for Theorem 3.3. This is the content of Theorem 3.7, where it turns out that, asymptotically, $L$ is either a local minimum or a saddle point, depending on its minimal vectors.
We complement the analysis of these two famous lattices with two instructive examples, illustrating how to carry out the analysis in very general cases.
In Section 5.4 we prove the following theorem.
Theorem 1.3.
There exists an extremal $2$-modular lattice in dimension $12$ which is not a critical point in the Gaussian core model.
This relates to a similar example in [14], where we illustrated this behavior at an even unimodular lattice in dimension $32$. What is different here is that the $32$-dimensional example came from a non-extremal lattice, so a lattice which cannot be optimal for sphere packing, while the new example is extremal. However, all $2$-modular lattices in dimension $12$ fail to be extreme (in the geometric sense), as they are neither perfect nor eutactic, which we checked by computer. In any case it is clear that the new $2$-modular example is not an optimal sphere packing, as it packs less dense than the $3$-modular Coxeter-Todd lattice.
We briefly mention that this being non-critical is not a property of all $2$-modular lattices in dimension $12$. Up to isometry there are $3$ such, and $2$ of them are non-critical, while one of them is. See the discussion in Section 5.4.
@@
Then, in Section 5.5, we illustrate how to deal with lattices for which the shells are spherical $2$-designs but not spherical $4$-designs. Then a more careful analysis is necessary, and this becomes rather technical if an associated space of cusp forms is of dimension at least $2$. This phenomenon occurs for extremal $\ell$-modular lattices, as opposed to the previous study of even unimodular lattices up to dimension $32$ in [14].
These techniques, together with explicit bounds on the coefficients of certain cusp forms, suffice to deal with all extremal $2$- and $3$-modular lattices in dimensions up to $20$ and $18$ respectively, these are the dimensions in which the classification of all extremal $2$- and $3$-modular lattices has been completed.
@@ 2.2. Theta series as modular forms
where $\Theta$ is the theta series of any even $\ell$-modular lattice of lowest possible dimension and $\Delta_{\ell}(\tau)=\left(\eta(\tau)\eta(\ell\tau)\right)^{24/(1+\ell)}$, where $\eta$ is the Dedekind eta function.
@@ 2.4. Extremal modular lattices and spherical designs
Let $L$ be an extremal $\ell$-modular lattice with $\ell$ prime and $1+\ell$ dividing $24$. Then we have vast knowledge about the minimal design strength of the shells of $L$.
This comes from a generalization of Venkov’s classical result regarding unimodular lattices [31]. Bachoc and Venkov collect this information in [1, Corollary 4.1]. We intend to use this for extremal $2$ and $3$-modular lattices.
@@
In all dimensions up to $18$ ($3$-modular) and $20$ ($2$-modular) all such lattices have been classified and additional data on them is readily available through the “Catalogue of lattices” [22]. We will collect some of this information in Section 5.1.
@@ 3.3. The Hessian in the presence of $2$-designs
Let $L$ be an even $\ell$-modular lattice. Let
@@ 3.4. Asymptotic behavior and properties of spaces of cusp forms
We frequently assume that $\mathcal{S}_{n/2+4}(\Gamma_{0}(\ell),\chi)$ has a basis in row echelon form. Whether such a basis exists, depending on the weight $n/2+4$, the prime $\ell$, and the character $\chi$ is, in general, unanswered. The existence of a basis in row-echelon form is guaranteed for the full modular group $\operatorname{SL}_{2}(\mathbb{Z})$, this is referred to as the (integral) Miller basis and is attributed to Miller [19]. However, even for $\Gamma_{0}(\ell)$, without a character, it can happen that such a basis no longer exists, for a recent treatment we refer to Wang [33]. In fact, the discussion after Question 1.5. in [33] gives that no “Miller basis”, i.e. basis in reduced row-echelon form, exists for $\mathcal{S}_{12k}(\Gamma_{0}(\ell))$ if $\ell$ is a prime and $312k<\ell-$ 3. The smallest such parameters give the space $\mathcal{S}_{12}(\Gamma_{0}(17))$, which is of dimension $16$ and admits a basis $C_{1},\ldots,C_{16}$ where the $q$-expansions of $C_{1},\ldots,C_{15}$ start with $q^{1},\ldots,q^{15}$, often followed by a multiple of $q^{16}$, but the $q$-expansion of $C_{16}$ starts with $q^{17}$.
@@
The first equality is contained in [27, Theorem 2.1], formulas for $\dim(\mathcal{S}_{n/2}(\Gamma_{*}(\ell),\chi))$ for certain $k$ and $\ell$ can be found in [23, Theorem 6.] and [27, Proposition 1.4]. We will be interested in the case that
@@
We briefly note that also the latter inequality in (26) can be verified directly, e.g. with the help of a computer, for $\ell$ and $k$ of interest. If $\chi$ acts trivially on $\Gamma_{0}(\ell)$ one can also easily reduce this question to evaluating the known dimension formulas for spaces of cusp forms for $\Gamma_{0}(\ell)$ (e.g. [12, Theorems 3.5.1 and 3.6.1]): In all such cases the dimension grows by at least $1$ since the dimension of $\mathcal{S}_{k}(\Gamma_{0}(\ell))$ grows at least with $\lfloor\frac{k}{4}\rfloor$.

### v2

@@ Abstract.
We discuss the local analysis of Gaussian potential energy of modular lattices. We present examples of $2$-modular lattices—such as the $16$-dimensional Barnes-Wall lattice—and $3$-modular lattices—such as the $12$-dimensional Coxeter-Todd lattice—that are locally universally optimal among lattices (in the sense of Cohn and Kumar). We also provide other $2$- and $3$-modular lattices that are not locally universally optimal, or not even critical in the Gaussian core model.
@@ 1.1. The Gaussian core model
for $\alpha>0$. Point configurations which interact via Gaussian potential functions $r\mapsto e^{-\alpha r^{2}}$ are referred to as the Gaussian core model. They are natural physical systems (see Stillinger [33]) and they are mathematically quite general. This embeds into the larger context of universal optimality: Following Cohn and Kumar [4], a point configuration is called universally optimal if it minimizes potential energy among all configurations of the same point density for every completely monotonic function of squared distance. It is an important observation that after fixing a point density, a point configuration is universally optimal if and only if it minimizes Gaussian potential energy for all $\alpha>0$. This is a consequence of a theorem of Bernstein (see Widder [38, Theorem 12b, page 161]). See Cohn, Kumar, Miller, Radchenko and Viazovska [6] for further discussion.
In the seminal paper [6] the authors proved that the $E_{8}$ root lattice in dimension $8$ and the Leech lattice $\Lambda_{24}$ in dimension $24$ are universally optimal point configurations in their respective dimensions. Further, the authors proved that these lattices are the unique minimizers among all periodic point configurations.
@@
One of the main goals of [16], motivated by Regev and Stephens-Davidowitz [29], was to find a concrete example of an even unimodular lattice that is a local maximum for some energy parameter $\alpha$ and the associated Gaussian potential. The first such example in this class of lattices can be found dimension $32$, every extremal even unimodular lattice in dimension $32$ is a local maximum for certain values of $\alpha$, a rigorous proof of this statement was worked out specifically for $\alpha=\pi$. It was proved that $E_{8}\perp E_{8}$ in dimension $16$ and every Niemeier lattice in dimension $24$ with a decomposable root system are saddle points if $\alpha$ is large enough. While all even unimodular lattices in dimensions up to $24$ turned out to be critical points in the Gaussian core model, there are non-critical even unimodular lattices in dimension $32$.
The aim of this paper is to extend the local analysis to a larger class of lattices, including dimensions in which no even unimodular lattice can exist. For this we consider the class of (extremal) modular lattices. This is in part motivated by an easy observation, which was already mentioned in Sarnak and Strömbergsson [30]: A lattice (of point density $1$) cannot be universally optimal if the theta series of $L$ and $L^{\#}$ are different. It seems noteworthy to mention that $A_{2}$ and $D_{4}$, which are locally universally optimal among lattices, are actually $2$-modular lattices, that is $A_{2}\cong\sqrt{2}A_{2}^{\#}$ and $D_{4}\cong\sqrt{2}D_{4}^{\#}$. From the general definition of an $\ell$-modular lattice below, it becomes clear that $\ell$-modular lattices satisfy this necessary condition and therefore seem good candidates to investigate if one is interested in universal optimality.
@@ 1.2. Modular lattices in the Gaussian core model
Modular lattices generalize unimodular lattices. This generalization is closely related to the theta series of a lattice $L$, which is a function of a variable $\tau$ in the complex upper half plane given by
@@
We now move to a larger class of lattices, for which $\Theta_{L}$ has similar desirable properties. Let $\ell>0$. A similarity $\sigma:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of norm $\ell$ is a linear map satisfying
@@
1.3. Structure of the paper
@@
The application of this strategy requires bounds on the coefficients of modular forms, which we obtain separately for Eisenstein series and cusp forms. This is the content of Section 4.
@@
1.4. Main results
We can summarize our results as follows. The main technical contribution of the paper is a technique that allows us to compute the eigenvalues of the Hessian of energy with respect to Gaussian potentials, if $L$ is an even $\ell$-modular lattice. This is the content of Theorem 3.3.
@@ 1.4. Main results
Building on this, if $L$ is an extremal even $\ell$-modular lattice, where $\ell$ is prime and $1+\ell$ divides $24$, we can determine the asymptotic behavior of $L$ in the Gaussian core model, as $\alpha\rightarrow\infty$, under the assumptions made for Theorem 3.3. This is the content of Theorem 3.7, where it turns out that, asymptotically, $L$ is either a local minimum or a saddle point, depending on its minimal vectors.
We apply the aforementioned techniques to $2$ and $3$-modular lattices in dimensions up to $20$ and $18$ respectively; these are the dimensions in which the classification of all extremal $2$- and $3$-modular lattices has been completed. We collect numerical and rigorous results on all of these cases in Tables 4 and 5, see Section 5.7 for more information on these tables.
@@
Specifically, while the tables contain a good amount of data which is only numerical evidence, this can be made rigorous. We illustrate this with an in depth example in Section 5.5: Here, we show how to deal with lattices for which the shells are spherical $2$-designs but not spherical $4$-designs. Then a more careful analysis is necessary, and this becomes rather technical if an associated space of cusp forms is of dimension at least $2$. This phenomenon occurs for extremal $\ell$-modular lattices, as opposed to the previous study of even unimodular lattices up to dimension $32$ in [16]. To illustrate the method, we apply it to one (out of six) extremal $3$-modular lattice in dimension $16$. Numerical evidence suggests that this lattice is a saddle point for all $\alpha$. We provide a rigorous proof of this for $\alpha=1$, to support this observation, and to illustrate the method. This is collected in Proposition 5.4.
In addition to the above, we discuss some cases where the shells of the lattices form stronger designs, namely at least $4$-designs. This includes the $3$-modular Coxeter-Todd lattice in dimension $12$ and the $2$-modular Barnes-Wall lattice in dimension $16$.
The main idea is to utilize that these lattices possess a lot of structure that simplifies the computation of the eigenvalues of the Hessian of $\mathcal{E}(\alpha,L)$ significantly; this comes from the fact that the shells of these lattices form spherical $4$-designs. Recall that if the first shell, i.e. the shell associated to minimal vectors of the lattice, forms a spherical $4$-design, then the lattice is called strongly perfect, a notion by Venkov [35]. It is known that strongly perfect lattices are extreme in the geometric sense, i.e. they are strict local maxima of the sphere packing density function for lattices. By a classical result of Voronoi [36] a lattice is extreme if and only if it is perfect and eutactic. Here a lattice $L$ is perfect if the span of the rank one matrices $xx^{\sf{T}}$, where $x$ runs through the set of minimal vectors of $L$, is equal to the space of symmetric matrices; $L$ is eutactic if the identity matrix is in the interior of the convex cone spanned by the rank one matrices $xx^{\sf{T}}$ from before.
We show that both lattices are locally universally optimal (see Propositions 5.1 and 5.2). This seems plausible as both lattices are the best known sphere packings in their respective dimension and the energy minimization problem converges to the sphere packing problem for $\alpha\rightarrow\infty$.
A direct extension of the proof of local universal optimality among lattices of these two well-known lattices, results in a criterion that asserts local universal optimality for other extremal $\ell$-modular lattices.
Theorem.
Let $L$ be an extremal $\ell$-modular lattice for which all shells are $4$-designs. Then, $L$ is locally universally optimal among lattices if $\ell\geq 3$, or if $\ell=2$ and $\dim(L)\in\left\{4,16,20,32,48\right\}$.
This theorem (Theorem 5.3) is proven in Section 5.4 and Table 3 collects all extremal $\ell$-modular lattices up to dimension $20$.
We complement the analysis in the presence of $2$-designs with a short proof (see Section 5.6) that there exists an extremal $2$-modular lattice in dimension $12$ which is not a universal critical point in the Gaussian core model.
What is different here is that the $32$-dimensional example came from a non-extremal lattice, so a lattice which cannot be optimal for sphere packing, while the new example is extremal. However, all $2$-modular lattices in dimension $12$ fail to be extreme (in the geometric sense), as they are neither perfect nor eutactic, which we checked by computer. In any case it is clear that the new $2$-modular example is not an optimal sphere packing, as it packs less dense than the $3$-modular Coxeter-Todd lattice.
We briefly mention that being non-critical is not a property of all $2$-modular lattices in dimension $12$. Up to isometry there are $3$ such, and $2$ of them are non-critical, while one of them is. See the discussion in Section 5.6 and Section 5.7.
@@ 2.2. Theta series as modular forms
where $\Theta$ is the theta series of any even $\ell$-modular lattice of lowest possible dimension and $\Delta_{\ell}(\tau)=\left(\eta(\tau)\eta(\ell\tau)\right)^{24/(1+\ell)}$, where $\eta$ is the Dedekind eta function. Here, the list of levels $\ell$ with $\ell+1\div 24$ is given by the primes $2,3,5,7,11,23$. Note that a similar algebra decomposition exists for a slightly larger class of levels, in particular the full level $\ell=1$, and the composite levels $6,14,15$ (see [28]).
@@ 2.4. Extremal modular lattices and spherical designs
There are two known, general ways to establish a minimum design strength on the shells of extremal $\ell$-modular lattices.
Firstly, Bachoc and Venkov [1] provide such bounds specifically for extremal $\ell$-modular lattices, when $\ell$ prime and $1+\ell$ divides $24$. This is a generalization of Venkov’s classical result regarding unimodular lattices [35], it can be found as Corollary 4.1 in [1]. We mainly use this for extremal $2$ and $3$-modular lattices.
@@
In all dimensions up to $18$ (33-modular) and 2020 (22-modular) all such lattices have been classified and additional data on them is readily available through the “Catalogue of lattices” [25]. We will collect some of this information in Section 5.1.
A second way, which works for arbitrary lattices, is based on a result of Goethals and Seidel [15]; we refer to this paper for all information exceeding the brief summary below. Let $L$ be a lattice, and consider the harmonic Molien series of $\operatorname{O}(L)$, the orthogonal group of $L$, which contains information about harmonic invariants of $\operatorname{O}(L)$: The $i$th coefficient of the harmonic Molien series of $\operatorname{O}(L)$ equals the dimension of the space of harmonic invariants of degree $i$ of $\operatorname{O}(L)$. Therefore, one can check that there are no $\operatorname{O}(L)$-invariant harmonic polynomials of degrees $1,\ldots,t$ for the isometry group $\operatorname{O}(L)$ of $L$, by computing the series and checking its coefficients; in practice we use the computer algebra system MAGMA [21] to do this for a specific lattice. This is helpful, since Theorem 3.12 in [15] asserts that all $\operatorname{O}(L)$-orbits are spherical $t$-designs if and only if there are no harmonic invariants of degree $1,\ldots,t$, this applies, in particular, to all shells of $L$, as they clearly are unions of orbits of $\operatorname{O}(L)$.
Note that this only provides a lower bound on the design strength of the shells of the lattice, as it considers all $\operatorname{O}(L)$-orbits, in fact one can find an example for this among the $2$-modular lattices in dimension $20$. There are $3$ such, and by the result of Bachoc and Venkov, all of their shells are at least $5$-designs. For two of these lattices this is also a consequence of the approach using the harmonic Molien series, but for one of them (the second entry in the file “2_dim20.dat” in [25]) the latter approach only yields a minimal strength of $3$, as there are harmonic invariants of degree $4$ for this specific lattice and its orthogonal group.
@@ 3.3. The Hessian in the presence of $2$-designs
Let $L$ be an even $\ell$-modular lattice, such that all shells of $L$ form spherical $2$-designs. Let
@@ 3.4. Asymptotic behavior and properties of spaces of cusp forms
We frequently assume that $\mathcal{S}_{n/2+4}(\Gamma_{0}(\ell),\chi)$ has a basis in row echelon form. Whether such a basis exists, depending on the weight $n/2+4$, the prime $\ell$, and the character $\chi$ is, in general, unanswered. The existence of a basis in row-echelon form is guaranteed for the full modular group $\operatorname{SL}_{2}(\mathbb{Z})$, this is referred to as the (integral) Miller basis and is attributed to Miller [22]. However, even for $\Gamma_{0}(\ell)$, without a character, it can happen that such a basis no longer exists, for a recent treatment we refer to Wang [37]. In fact, the discussion after Question 1.5. in [37] gives that no “Miller basis”, i.e. basis in reduced row-echelon form, exists for $\mathcal{S}_{12k}(\Gamma_{0}(\ell))$ if $\ell$ is a prime and $12k<\ell-3$. The smallest such parameters give the space $\mathcal{S}_{12}(\Gamma_{0}(17))$, which is of dimension $16$ and admits a basis $C_{1},\ldots,C_{16}$ where the $q$-expansions of $C_{1},\ldots,C_{15}$ start with $q^{1},\ldots,q^{15}$, often followed by a multiple of $q^{16}$, but the $q$-expansion of $C_{16}$ starts with $q^{17}$.
@@
The first equality is contained in [31, Theorem 2.1], formulas for $\dim(\mathcal{S}_{n/2}(\Gamma_{*}(\ell),\chi))$ for certain $k$ and $\ell$ can be found in [27, Theorem 6.] and [31, Proposition 1.4]. We will be interested in the case that
@@
We briefly note that also the latter inequality in (26) can be verified directly, e.g. with the help of a computer, for $\ell$ and $k$ of interest. If $\chi$ acts trivially on $\Gamma_{0}(\ell)$ one can also easily reduce this question to evaluating the known dimension formulas for spaces of cusp forms for $\Gamma_{0}(\ell)$ (e.g. [14, Theorems 3.5.1 and 3.6.1]): In all such cases the dimension grows by at least $1$ since the dimension of $\mathcal{S}_{k}(\Gamma_{0}(\ell))$ grows at least with $\lfloor\frac{k}{4}\rfloor$.